# Wildfire Risk Prediction: A Review

### v3

@@ Abstract
Wildfires significantly impact global ecosystems, wildlife, and human populations, destroying plant communities, habitats, and settlements. They also contribute to increased emissions of pollutants such as carbon dioxide, nitrogen oxides, and methane. Thus, accurate wildfire risk prediction is essential for reducing emissions, mitigating pollution, and protecting biodiversity and human health. However, predicting wildfire risk is a complex challenge that involves defining risk, integrating multi-source data, selecting appropriate algorithms, ensuring algorithm interpretability, and calibrating predictions.
This paper addresses the entire wildfire risk prediction process. It first introduces the concept of wildfire risk and summarizes the geographic distribution of referenced studies. It then reviews commonly used factors in wildfire risk prediction, such as fuel, meteorological, and socio-economic data, along with associated sources and computational methods. Additionally, we explore the issue of feature collinearity in data, presenting approaches to evaluate collinearity and attribute feature importance, with an emphasis on addressing the black-box nature of machine learning models.
We further provide an overview of wildfire danger rating systems and traditional machine learning algorithms. Given the rapid advancement of deep learning and its increasing application in wildfire risk assessment, we identify a gap in the systematic review of these methods. To bridge this gap, we categorize and review deep learning algorithms into three groups: time series prediction, image segmentation and classification, and spatiotemporal prediction. We also discuss methods for converting model outputs into actionable probabilities. Finally, we offer insights and propose future directions for deep learning-based wildfire prediction.
@@
wildfire prediction , remote sensing , machine learning , review
@@ 1 Introduction
Wildfires occur in diverse environments such as forests, grasslands, farmlands, and peatlands, with significant impacts on global ecosystems, wildlife, human health, air quality, and carbon balance. They not only destroy plant communities, habitats, and human settlements but also contribute to elevated emissions of greenhouse gases and toxic pollutants like carbon dioxide, nitrogen oxides, methane, ozone, particulate matter, and mercury (Pelletier et al., 2023; Hua and Shao, 2017; McNorton and Di Giuseppe, 2024; Yang et al., 2022). For instance, the devastating 2019 Australian wildfires burned approximately 18.6 million hectares of land, destroyed over 5,900 structures, and caused the deaths of at least 34 people and 1 billion animals (Naderpour et al., 2020). In the coming decades, climate change is expected to shorten fire return intervals and increase the burned area in vulnerable regions (Parks and Abatzoglou, 2020; Holdrege et al., 2024). As a result, accurate wildfire risk prediction is essential for mitigating future impacts.
@@
In response to these challenges, wildfire risk prediction has attracted growing attention from governments, academia, and industry. Current prediction models consider essential components such as defining the study area and timescale, data selection and processing, method selection, setting prediction objectives, and evaluating model accuracy. However, a systematic review of these elements remains lacking. This paper addresses this gap by providing a comprehensive review of the data, workflows, and prediction methods used for wildfire risk assessment, while also outlining potential future research directions.
@@
A major challenge in wildfire risk prediction lies in the lack of a clear and consistent definition of risk across studies. Based on the literature, this paper categorizes wildfire risk into four types: explicit definitions, implicit definitions, ignition risk, and burning risk. Wildfire ignition and spread result from complex interactions among various factors, including fuel conditions, climate, weather, socio-economic factors, terrain, hydrology, and historical wildfire data. This paper details the definitions, calculations, proxies, and frequency of use of these features, alongside an overview of publicly available spatial datasets for wildfire spread prediction.
@@
A key concern in the design of prediction algorithms is evaluating feature collinearity and improving model interpretability. Including highly correlated features can degrade model performance and obscure feature importance, while model interpretability helps to clarify patterns of ignition and spread. This paper introduces techniques for assessing feature collinearity and enhancing the interpretability of machine learning models.
@@
Another critical aspect of wildfire prediction is the choice of algorithms. Methods range from wildfire danger rating systems and statistical models to traditional machine learning and deep learning approaches. While Jain et al. (2020) provides a comprehensive review of machine learning models, this paper focuses on deep learning, which has seen rapid adoption and is becoming a crucial tool for wildfire prediction. To address the lack of systematic reviews in this area, we categorize deep learning-based models into three task types: time series prediction, image segmentation and classification, and spatiotemporal prediction. This study reviews existing deep learning models in these categories, providing a structured overview of their applications.
@@
Finally, with the ongoing advancements in satellite observation, integrated air-ground-space remote sensing technologies, and deep learning models, wildfire risk prediction using multi-source remote sensing presents significant opportunities and challenges. This paper concludes by discussing the current limitations in data and algorithms and highlighting potential future developments.
@@ 2 Wildfire risk definition
Figure 1: The generalized wildfire risk framwork proposed by Miller and Ager (2012)
Additionally, their study distinguishes between commonly used terms in wildfire risk research, such as risk, hazard, exposure, threats, vulnerability, and fire danger. It is important to note that most of the studies summarized in this paper focus more on the threats posed by wildfires to human life and property rather than solely on environmental and ecological impacts. Therefore, these studies are often conducted in areas with high social value density. Based on the aforementioned definitions, wildfire risk can be approximately equated to wildfire likelihood.
When summarizing related work, we focus only on studies concerning wildfire risk and danger, excluding those on hazard, exposure, and threats. The former primarily describes the outlook for fire occurrence over a certain period, while the latter tends to evaluate the hazard of wildfires (Fig.1). Consequently, if we consider only the likelihood of fire occurrence without considering its potential harm, the definition of wildfire risk can be simplified to the probability of a wildfire igniting or spreading at a specific location and time (Deeming, 1972). Many researchers in the past have described risk as the likelihood of wildfire occurrence, which also aligns with the wildfire framework established by the European Commission (Depicker et al., 2020).
@@
2.1 Explicit definition methods
This section focuses on explicit definition methods, which include directly using historical observation data to define wildfire risk, defining wildfire risk based on the weights of influencing factors, and combining these factors to simulate wildfire distribution and define risk through simulation results. It is worth noting that this subsection does not cover the details of the simulation algorithms, but rather how to convert simulation results into wildfire risk.
Firstly, the simplest and most direct method in wildfire risk estimation is using historical fire occurrences to predict future risk. For example, Rubí and Gondim (2023) classified regions where one or more fires occurred within an hour as "high risk," while other areas were classified as "low risk." Preisler et al. (2004) selected all fire voxels within a past time ($t$) and spatial range and sampled a $\pi$ percentage of non-fire voxels using the scheme proposed by Brillinger et al. (2003). The occurrence of fire can then be characterized using conditional probability, which can be constrained within the [0, 1] interval using a sigmoid function.
@@ 2.1 Explicit definition methods
where $m_{i,j}$ represents the fire ignition probability in grid cell $g$ during period $j$; $n_{g,j}$ and $N$ are the total number of wildfires and the total number of considered periods, respectively. They also introduced Monte Carlo techniques to interpolate grid cells. Xu et al. (2023) modeled wildfire risk using a marked spatiotemporal Hawkes process, with the conditional intensity function before location $k$ and time $t$.
$\lambda(t,k,m|\mathcal{H}_{t})=$ (3)
$\lim_{\Delta t,\Delta u\to 0}\frac{\mathbb{E}[\mathbb{N}(|t,t+\Delta t)\times B(k,\Delta k)\times B(m,\Delta m))|\mathcal{H_{t}}]}{\Delta\times B(k,\Delta k)\times B(m,\Delta m)}$
where $B(a,r)$ is a sphere with center $a$ and radius $r$, and $\mathcal{N}$ is the counting measure. Lall and Mathibela (2016) assessed wildfire risk in Cape Town, South Africa, by defining daily risk ratings based on the average number of wildfires per day and their corresponding standard deviation over seven years. Specifically, fire occurrences below the average were rated as 1 or lower, those between the average and one standard deviation above were rated as 2 or moderate, those between one and two standard deviations above were rated as 3 or higher, and occurrences more than two standard deviations above the average were rated as 4 or extreme. An ANN model was then established using temperature, humidity, wind speed, and precipitation as inputs to predict wildfire risk.
The above models based on historical data do not consider the conditions for wildfire ignition or spread, thus having poor robustness and generalization ability. Moreover, due to the infrequent occurrence of wildfires over long time scales and large areas, there is a severe class imbalance between the occurrence and non-occurrence of wildfires in space and time. Therefore, the risk assessment results of these models are greatly influenced by data sampling. Consequently, some studies have begun to incorporate or solely use factors such as fuel, terrain, weather conditions, and wildfire triggers as proxies to estimate ignition risk. For example, Castro et al. (2003) used shrub moisture content, specifically live fuel moisture, as a proxy variable for wildfire risk in Spain. More complexly, Sadasivuni et al. (2013) defined wildfire risk as the interaction between human activity and fuel distribution. Specifically, they integrated population and fuel layers through additive operations as input variables for a gravity model in a spatial interaction model. The prediction results were then classified into five levels. Similarly, Gentilucci et al. (2024) categorized wildfire influencing factors and input them into a weight of evidence model, categorizing wildfire risk into low and high levels through regression of various influencing factor weights. Oliveira et al. (2021a); Bergonse et al. (2021) conducted comprehensive classifications of land use, land cover, slope, elevation, and aspect. They then calculated the likelihood ratio for each category and determined fire susceptibility by summing the weight values of different categories. This process can be mathematically represented as follows:
@@
Srivastava et al. (2014) differed from the aforementioned studies by simultaneously considering both causal and anti-causal factors through an expert knowledge system to quantify the likelihood of forest fires.
@@
Wildfire ignition and spread are highly random and exhibit a strong nonlinear relationship with various factors. Data-driven machine learning methods are often used in wildfire risk prediction research due to their ability to fit nonlinear functions effectively with large-scale data. However, these methods cannot explicitly express the influence of various factors on the final wildfire risk. Specifically, commonly used machine learning-based wildfire risk probabilities use binary historical wildfire event datasets as labels, with multiple features as input parameters, defining the binary output or probability output as wildfire risk.
For example, Yahia et al. (2023) used the normalized binary burn ratio as a proxy for wildfire risk probability. This probability, along with proxy variables such as temperature vegetation dryness, vertical drought index, differential water index, and indices derived from the optical trapezoid model, were used to train a Gaussian naive Bayes classifier. Malik et al. (2021) similarly used historical binary wildfire occurrence records to train an ensemble model combining Adaboost and two random forest models to predict wildfire risk probability. Qiao et al. (2024) constructed a Transformer-based deep learning wildfire risk prediction model, using the output of an MLP as the predicted risk probability. Similarly, Ghorbanzadeh et al. (2019) used MODIS historical wildfire records to train a multilayer perceptron model, with the logistic function output of the model representing wildfire susceptibility. Jiang et al. (2024b) also used MODIS historical wildfire events as true values but created a balanced dataset by selecting non-fire areas in a 1:1 ratio to fire points and used a deep learning model to predict risk probability maps. Botequim et al. (2013) used annual wildfire occurrence records as the dependent variable to regress a logistic function and plotted wildfire occurrence probability. They then used a sigmoid function to describe the relationship between the independent and dependent variables:
$y=\frac{1}{1+e^{-(\beta_{0}+\beta_{1}x_{1}+...+\beta_{n}x_{n})}}$ (6)
where $x_{i}$ and $y$ represent the $i$-th independent variable and historical wildfire occurrence records, respectively. After calculating the regression coefficients $\beta_{i}$, the logistic model uses the independent variables to output a continuous wildfire occurrence probability between 0 and 1.
@@
It is important to note that the wildfire risk defined by the aforementioned machine learning models is essentially a pseudo-probability rather than a true probability density (Pelletier et al., 2023). For example, the predicted pseudo-probability in XGBoost is the inverse log transformation of the sum of terminal leaf weights; in SVM, it is estimated by calculating the distance to the classification decision boundary and using these distances for probability estimation; in Random Forest, it is the average of the voting results from multiple decision trees; and in Multilayer Perceptron (MLP), Long Short-Term Memory (LSTM), and Convolutional Neural Network (CNN), it is typically the probability estimate from the final layer activation function. Furthermore, compared to explicit definitions, these models do not provide specific impacts of various features on wildfire occurrence and strongly depend on the construction of the label sample set and the choice of model.
Specifically, different preprocessing methods for historical wildfire data, as well as model selection and performance, can greatly influence the predicted probability values. For instance, when using wildfire influencing factors and historical records at time $t$ as model inputs and selecting historical records at time $t+1$ as labels to train the model, the model, if performing well, will predict high probabilities for positive regions at time $t+1$, such as greater than 0.8. However, wildfire risk probabilities are usually low; for example, Ager et al. (2014a) and Preisler et al. (2004) used millesimal probabilities in their wildfire risk assessments for the Mediterranean region and the United States, respectively. Therefore, it is necessary to perform post-processing on the probability predictions, such as establishing a link between explicit and implicit definitions.
2.2 Wildfire ignition risk
@@
It is important to note that the studies mentioned above utilized ground-based wildfire databases rather than the more commonly used MODIS or VIIRS fire/hotspot data in wildfire risk assessment research. This may be due to the relatively low spatiotemporal resolution of these data, making it difficult to detect ignition points with very low intensity and duration. Benali et al. (2016) compared the consistency of wildfire field records in Portugal, Greece, Alaska, California, and southeastern Australia with MODIS active fire products (MCD14ML). They proposed a clustering algorithm to calculate the ignition and end dates of a wildfire in the MCD14ML product and used the Nash–Sutcliffe Model Efficiency Index to estimate the consistency between satellite-estimated fire dates and ground records. The study found that satellite data could not estimate the start and end dates of most fires smaller than 500 ha. Similarly, Coskuner (2022) compared the performance of MODIS and S-NPP VIIRS active fire or hotspot detection products, i.e., MCD14ML V6 and VNP14IMG, in detecting wildfires across five different land cover types in Turkey and compared these detections with records from Turkey’s ground-based wildfire database. The authors found that both MODIS and VIIRS had very low detection capabilities for small fires (<1 ha) (<3.5%). They also cited studies by Fusco et al. (2019) and Ying et al. (2019), showing that MODIS fire products had high omission rates in the United States and China. Specifically, Fusco et al. (2019) reported detection rates for MODIS products in the United States ranging from 3.5% to 23.4%, with detection rates increasing with fire size. Similarly, Ying et al. (2019) found that in Yunnan Province, China, only 11.10% of 5,145 ground-recorded wildfire events were detected by MODIS fire products, with most omissions attributed to the small size of the fires. These analyses suggest that the commonly used MODIS and VIIRS fire and hotspot products are not suitable for estimating ignition locations and events, nor for corresponding ignition risk estimation.

### v4

# Wildfire Risk Prediction: A Survey of Recent Advances Using Deep Learning Techniques

@@ Abstract
Wildfires pose significant threats to ecosystems, wildlife, and human populations, leading to habitat destruction, pollutant emissions, and biodiversity loss. Accurate wildfire risk prediction is crucial for mitigating these impacts and protecting both environmental and human health. This paper presents a comprehensive review of wildfire risk prediction methodologies, particularly focusing on deep learning approaches. It begins by defining wildfire risk and summarizing the geographical distribution of relevant studies. Key predictive features, including meteorological, socio-economic, and fuel-related factors, are analyzed, with emphasis on feature collinearity evaluation and model interpretability to enhance predictive accuracy. Traditional methods such as wildfire danger rating systems and machine learning models are reviewed, followed by a categorization of deep learning models into time series prediction, image segmentation, and spatiotemporal prediction. Methods for translating model outputs into actionable probabilities are also discussed. The paper identifies the current challenges and limitations in existing models and outlines future opportunities for multimodal data integration and the development of architecturally efficient models for improved wildfire risk prediction.
@@
wildfire , risk prediction , remote sensing , deep learning , review
@@ 1 Introduction
Wildfires occur in diverse environments such as forests, grasslands, farmlands, and peatlands, with significant impacts on global ecosystems, wildlife, human health, air quality, and carbon balance (Moritz et al., 2014; Stojanovic et al., 2016; Bowd et al., 2021; Holder et al., 2023; Lopez et al., 2024). They not only destroy plant communities, habitats, and human settlements but also contribute to elevated emissions of greenhouse gases and toxic pollutants like carbon dioxide, nitrogen oxides, methane, ozone, particulate matter, and mercury (Hua and Shao, 2017; Yang et al., 2022a; Pelletier et al., 2023; McNorton and Di Giuseppe, 2024). For instance, during the record-breaking 2023 wildfire season, approximately 18.5 million hectares burned across Canada (Erni et al., 2024), the largest area recorded since national reporting began in 1972. This season resulted in the evacuation of 232,000 people (Byrne et al., 2024), with over 200 communities evacuated, hundreds of homes, and infrastructure destroyed (Jain et al., 2024). The carbon emissions for this season were estimated at 647 Tg C (Byrne et al., 2024), with potential long-term impacts on communities and ecosystems expected to last for years or even decades. In the coming decades, climate change is expected to shorten fire return intervals and increase the burned area in vulnerable regions (Parks and Abatzoglou, 2020; Holdrege et al., 2024).
@@
In response to these challenges, wildfire risk prediction has attracted growing attention from governments, academia, and industry. Current prediction models consider essential components such as research scope delineation, data selection and processing, method establishment, setting prediction objectives, and evaluating model accuracy. However, a systematic review of these elements remains lacking. This paper addresses this gap by providing a comprehensive review of the data and deep learning methods used for wildfire risk assessment. In addition, the popular feature attribution techniques in deep-leaning-based wildfire risk prediction models are also reviewed. Finally, we also outline promising new deep learning methods.
@@
First, wildfire ignition and spread result from complex interactions among various factors, including fuel conditions, climate, weather, socio-economic factors, terrain, hydrology, and historical wildfire data (Fuller, 1991). This paper details the definitions, calculations, proxies, and frequency of use of these features, alongside an overview of publicly available spatial datasets for wildfire spread prediction.
@@
Second, A key concern in the design of prediction algorithms is evaluating feature collinearity and improving model interpretability. Including highly correlated features can degrade model performance and obscure feature importance (Li et al., 2022a), while model interpretability helps to clarify patterns of ignition and spread (Abdollahi and Pradhan, 2023). This paper introduces techniques for assessing feature collinearity and enhancing the interpretability of machine learning models.
@@
Another critical aspect of wildfire prediction is the algorithm establishment. Methods range from wildfire danger rating systems and statistical models to traditional machine learning and deep learning approaches. While Jain et al. (2020) provides a comprehensive review of machine learning models, this paper focuses on deep learning, which has seen rapid adoption and is becoming a crucial tool for wildfire predictions. To address the lack of systematic reviews in this area, we categorize deep learning-based models into three task types: time series prediction, image segmentation and classification, and spatiotemporal prediction. This study reviews existing deep learning models in these categories, providing a structured overview of their applications.
@@
Finally, with the continuous advancements in deep learning techniques, the integration of multi-source geospatial data and deep learning models for wildfire risk prediction offers both significant opportunities and challenges. This paper also discusses the current limitations of existing algorithms and highlighting potential future developments in multimodal and architecturally efficient models.
@@ 2 Wildfire Risk Definition
Figure 1: The generalized wildfire risk framework proposed by Miller and Ager (2012)
Additionally, their study distinguishes between commonly used terms in wildfire risk research, such as risk, hazard, exposure, threats, vulnerability, and fire danger. When summarizing related work, we focus only on studies concerning wildfire risk and danger, excluding those on hazard, exposure, and threats. The former primarily describes the outlook for fire occurrence over a certain period, while the latter tends to evaluate the hazard of wildfires (Fig.1). Consequently, if we consider only the likelihood of fire occurrence without considering its potential harm, the definition of wildfire risk can be simplified to the probability of a wildfire igniting or spreading at a specific location and time (Deeming, 1972). Many researchers in the past have described risk as the likelihood of wildfire occurrence, which also aligns with the wildfire framework established by the European Commission (Depicker et al., 2020).
@@
2.1 Explicit and Implicit Definition Methods
This subsection focuses on explicit definition methods, which include directly using historical observation data to define wildfire risk, defining wildfire risk based on the weights of influencing factors, and combining these factors to simulate wildfire distribution and define risk through simulation results. It is worth noting that this subsection does not cover the details of the simulation algorithms, but rather how to convert simulation results into wildfire risk.
Firstly, the simplest and most direct method in wildfire risk estimation is using historical fire occurrences to predict future risk. For example, Rubí and Gondim (2023) classified regions where one or more fires occurred within an hour as "high risk," while other areas were classified as "low risk." Preisler et al. (2004) selected all fire voxels within a past time ($t$) and spatial range and sampled a $\pi$ percentage of non-fire voxels using the scheme proposed by Brillinger et al. (2003). The occurrence of fire can then be characterized using conditional probability, which can be constrained within the (0, 1) interval using a sigmoid function.
@@ 2.1 Explicit and Implicit Definition Methods
where $m_{g,j}$ represents the fire ignition probability in grid cell $g$ during period $j$; $n_{g,j}$ and $N$ are the total number of wildfires and the total number of considered periods, respectively. They also introduced Monte Carlo techniques to interpolate grid cells. Lall and Mathibela (2016) assessed wildfire risk in Cape Town, South Africa, by defining daily risk ratings based on the average number of wildfires per day and their corresponding standard deviation over seven years. Specifically, fire occurrences below the average were rated as 1 or lower, those between the average and one standard deviation above were rated as 2 or moderate, those between one and two standard deviations above were rated as 3 or higher, and occurrences more than two standard deviations above the average were rated as 4 or extreme. An ANN model was then established using temperature, humidity, wind speed, and precipitation as inputs to predict wildfire risk.
The above models based on historical data do not consider the conditions for wildfire ignition or spread, thus having poor robustness and generalization ability. Moreover, due to the infrequent occurrence of wildfires over long time scales and large areas, there is a severe class imbalance between the occurrence and non-occurrence of wildfires in space and time. Therefore, the risk assessment results of these models are greatly influenced by data sampling. Consequently, some studies have begun to incorporate or solely use factors such as fuel, terrain, weather conditions, and wildfire triggers as proxies to estimate ignition risk. For example, Castro et al. (2003) used shrub moisture content, specifically live fuel moisture, as a proxy variable for wildfire risk in Spain. More complexly, Sadasivuni et al. (2013) defined wildfire risk as the interaction between human activity and fuel distribution. Specifically, they integrated population and fuel layers through additive operations as input variables for a gravity model in a spatial interaction model. The prediction results were then classified into five levels. Similarly, Gentilucci et al. (2024) categorized wildfire influencing factors and input them into a weight of evidence model, categorizing wildfire risk into low and high levels through regression of various influencing factor weights. Oliveira et al. (2021a); Bergonse et al. (2021) conducted classifications of land use, land cover, slope, elevation, and aspect. They then calculated the likelihood ratio for each category and determined fire susceptibility by summing the weight values of different categories. This process can be mathematically represented as follows:
@@
where FDI represents the fire danger index. Srivastava et al. (2014) differed from the aforementioned studies by simultaneously considering both causal and anti-causal factors, including the distribution of check dams, firelines, and fire water towers, using an expert knowledge system to quantify the likelihood of forest fires.
@@
On the other hand, wildfire ignition and spread are highly random and exhibit a strong nonlinear relationship with various factors (Leuenberger et al., 2018). Data-driven machine learning methods are frequently employed in wildfire risk prediction research due to their capability to effectively model nonlinear functions using large-scale datasets (Jain et al., 2020). However, these methods do not explicitly quantify the contribution of individual factors to the overall wildfire risk, compared to explici definition methods. Thus, these methods can be categorized as implicit definition methods. Further details on these methods will be provided in Section 7.
@@
It is important to note that the wildfire risk defined by these machine learning models represents a pseudo-probability rather than a true probability density (Pelletier et al., 2023). This pseudo-probability arises from the models’ limitations in explicitly defining risk. For example, in XGBoost, the predicted pseudo-probability is derived from the inverse log transformation of the sum of terminal leaf weights. In SVM, it is estimated by calculating the distance to the classification decision boundary and using these distances for probability estimation. In Random Forest, it is the average of the voting outcomes from multiple decision trees. For models such as Multilayer Perceptron (MLP), Long Short-Term Memory (LSTM), and Convolutional Neural Network (CNN), it is typically the probability estimate from the activation function of the final layer.
Apart from the impact of method selection on the prediction probability, these models also heavily depend on the construction of the label sample set. Different preprocessing techniques, such as historical resampling of wildfire data, can significantly affect the predicted probability values. Therefore, it is necessary to apply post-processing techniques, such as establishing a connection between explicit and implicit definitions, to enhance the interpretability and accuracy of the model outputs.
2.2 Wildfire Ignition and Burring Risk
@@
It is important to note that the studies mentioned above utilized ground-based wildfire databases rather than the more commonly used MODIS or VIIRS fire/hotspot data in wildfire risk prediction research. This may be due to the relatively low spatiotemporal resolution of these data, making it difficult to detect ignition points with very low intensity and duration. Benali et al. (2016) compared the consistency of wildfire field records in Portugal, Greece, Alaska, California, and southeastern Australia with MODIS active fire products (MCD14ML). They proposed a clustering algorithm to calculate the ignition and end dates of a wildfire in the MCD14ML product and used the Nash–Sutcliffe Model Efficiency Index to estimate the consistency between satellite-estimated fire dates and ground records. The study found that satellite data could not estimate the start and end dates of most fires smaller than 500 ha. Similarly, Coskuner (2022) compared the performance of MODIS and S-NPP VIIRS active fire or hotspot detection products, i.e., MCD14ML V6 and VNP14IMG, in detecting wildfires across five different land cover types in Turkey and compared these detections with records from Turkey’s ground-based wildfire database. The authors found that both MODIS and VIIRS had very low detection capabilities for small fires (<1 ha) (<3.5%). They also cited studies by Fusco et al. (2019) and Ying et al. (2019), showing that MODIS fire products had high omission rates in the United States and China. Specifically, Fusco et al. (2019) reported detection rates for MODIS products in the United States ranging from 3.5% to 23.4%, with detection rates increasing with fire size. Similarly, Ying et al. (2019) found that in Yunnan Province, China, only 11.10% of 5,145 ground-recorded wildfire events were detected by MODIS fire products, with most omissions attributed to the small size of the fires. These analyses suggest that the commonly used MODIS and VIIRS fire and hotspot products are not suitable for estimating ignition locations and events, nor for corresponding ignition risk estimation.